OpenStudy (anonymous):

simplify completely 4m^3/3m^2n multiplied by 9n^2/30m^2n^2

OpenStudy (anonymous):

2/(5mn)

OpenStudy (anonymous):

\[\frac{4m^3}{2m^2n}=\frac{4m^{3-2}}{3n}=\frac{4m}{3n}\] is a start

OpenStudy (anonymous):

\[\frac{9n^2}{30m^2n^2}\] \[=\frac{3}{10m^2}\]

OpenStudy (anonymous):

multiply them together get \[\frac{4m}{3n}\times \frac{3}{10m^2}\] \[=\frac{4}{10mn}=\frac{2}{5mn}\]

OpenStudy (anonymous):

Thank you!!!